# Online Intrinsic Rewards for Decision Making Agents from Large Language Model Feedback

### v3

@@ 5.2 Comparisons for More Complex Goals
Even though ELLM-BoW performs well in Section 5.1, we have found that it does not capture the semantic meanings of more complex goal strings due to the simple bag-of-word representation. To demonstrate this, we train ELLM-BoW and ONI-retrieval for two opposite goals in the extrinsic-reward-free environment: (Gold) “collect gold but do not kill monsters” vs (Combat) “kill monsters but do not collect gold”. Figure 4 shows that ELLM-BoW produces two agents of nearly identical behavior in terms of collected gold and killed monsters. In contrast, ONI-retrieval is able to distinguish the two goals and produces agents that emphasize either combat engagement or gold collection, depending on the goal string. See Section B.1 for the goal strings for ELLM-BoW and prompts for ONI-retrieval.
@@ 5.3 Ablation Studies
Performance vs. LLM Annotation Throughput We compared agents trained on the Score and Oracle tasks using either $1$ or $4$ V100 GPUs in the LLM server node. As shown in Figure 5, using 4 GPUs rather than 1 significantly increases the number of annotated observations, but does not significantly change the performance. This suggests that many of the labelled examples may contain redundant information, which is not useful for updating the reward model. Designing more sophisticated prioritization schemes, which can select maximally informative examples to send to the LLM, constitutes an interesting direction for future work.
@@
Performance when Reducing LLM Annotations Thus far we have been using all LLM annotations available. It is also intriguing to check the performance when we limit the volume of annotations, to simulate more resource-constrained settings. Here we subsample the LLM-annotated messages with rate 0.1 and 0.01 before sending them back to the hash table. Figure 6 shows that the performance of ONI-retrieval significantly drops with rate 0.01, whereas ONI-classification remains comparable. This suggests that using a parametric reward model can help reduce the number of annotations required thanks to its generalization ability.
@@ A.2 Hyperparameters
Table 1 shows the APPO hyperparameters which are common to all experiments and Table 2 includes the hyperparameters for the online LLM annotation.
@@ Appendix B Prompts
B.1 Goal String and Prompts used in Section 5.2
@@ B.1.2 Prompt used for ONI-retrieval
The prompt used for ONI methods have been modified to reflect the change of goals, which we list in the next subsection. In addition to the changes of goal description, we use the same examples as in Appendix B but changed its label.
@@ Appendix C Additional Results
In Figure 7, we compare the performance of Motif using two different sized LLMs on Score and Oracle (LLaMA-3.1-8B-Instruct and LLaMA-3.1-70B-Instruct). Interestingly, we do not observe a significant difference between the two. This is in contrast to the previous work of (Klissarov et al., 2023), who found a significant difference between using LLaMA-2-70B-chat and LLaMA-2-7B-chat. This suggest that the smaller 8B model is sufficient for evaluating messages on these tasks, hence we use it in our experiments.
Figure 7: Performance of Motif using two different LLMs. Curves represent means and shaded regions represent standard errors over $5$ seeds.
@@ C.2 Effect of the Episodic Term for ELLM-BoW
Our implementation of ELLM-BoW has included the episodic-count based normalization (6), which is key to the performance of ELLM-BoW. Figure 8 shows that directly using ELLM-BoW as in Du et al. (2023a) ELLM-BoW, without the episodic term, failed to make progress in all three tasks.
@@ C.3 Additional Ablations
Impact of Intrinsic Reward Coefficient $\beta$ For ONI-classification and ONI-retrieval, we have used different values of $\beta$ for the Score task and other sparse reward tasks in Section 5.1. Throughout our experiments, we have found that for these two methods, larger values of $\beta$ lead to better performance for the sparse reward tasks, while smaller values of $\beta$ better balance between intrinsic and extrinsic rewards for the dense reward Score task. In comparison, ONI-ranking is relatively robust to this choice, where larger values of $\beta$ are still slightly favored for the sparse reward tasks.
ONI-classification: the impact of the classification threshold As described in Section 3.2, ONI-classification predicts binary rewards by modeling $P(y_{t}=1|o_{t})$ and then thresholding with $\eta$. Instead of using binary reward, an alternative design choice is to use real valued reward
@@
where $r^{\mathrm{int}}(o_{t})\in[0,1]$ is the output of the reward classifier before thresholding. Figure 11 shows that the performance of ONI-classification on the four NetHack tasks is relatively robust to this hyperparameter $\eta$, and using $P(y_{t}=1|o_{t})$ as the reward leads to similar performance. As the training progresses, our reward model outputs values close to 0 or 1777To increase the diversity of the captions in the training dataset, we do not requery the LLM server if a caption is already annotated before. Therefore, every caption in our dataset only has a single label, resulting in this phenomenon., and we hypothesize this is the reason why the final performance remains comparable. Moreover, the natural classification threshold $\eta=0.5$ marginally outperforms the other values in the reward-free setting.
@@
Figure 12 shows the performance and the number of captions stored in the replay buffer for both schemes. We see that for the deduplicated variant, the number of captions does not grow past a certain point (approximately $70k$ unique captions). For the non-deduplicated variant, the number of captions keeps growing linearly over time. The deduplicated variant fails to learn, which highlights the important effect which the annotation dataset can have. We hypothesize that the deduplicated variant may undersample captions that occur frequently in the agent’s experience, for which it is important to reliably estimate reward. For example, the blank message occurs very frequently during policy learning, but is only included in a small fraction of the pairs sent to the LLM for annotation, since it is sampled with the same probability as the other $\sim 70k$ other captions. In contrast, the duplicated variant samples the blank message and other frequent captions with much higher probability. Still, this remains a simple strategy, and designing more sophisticated sampling mechanisms (for example, that account for the epistemic uncertainty of the reward model) would be an interesting direction for future work.

### v4

@@ 5.2 Comparisons for More Complex Goals
Even though ELLM-BoW performs well in Section˜5.1, we have found that it does not capture the semantic meanings of more complex goal strings due to the simple bag-of-word representation. To demonstrate this, we train ELLM-BoW and ONI-retrieval for two opposite goals in the extrinsic-reward-free environment: (Gold) “collect gold but do not kill monsters” vs (Combat) “kill monsters but do not collect gold”. Figure˜4 shows that ELLM-BoW produces two agents of nearly identical behavior in terms of collected gold and killed monsters. In contrast, ONI-retrieval is able to distinguish the two goals and produces agents that emphasize either combat engagement or gold collection, depending on the goal string. See Section˜B.1 for the goal strings for ELLM-BoW and prompts for ONI-retrieval.
@@ 5.3 Ablation Studies
Performance vs. LLM Annotation Throughput We compared agents trained on the Score and Oracle tasks using either $1$ or $4$ V100 GPUs in the LLM server node. As shown in Figure˜5, using 4 GPUs rather than 1 significantly increases the number of annotated observations, but does not significantly change the performance. This suggests that many of the labelled examples may contain redundant information, which is not useful for updating the reward model. Designing more sophisticated prioritization schemes, which can select maximally informative examples to send to the LLM, constitutes an interesting direction for future work.
@@
Performance when Reducing LLM Annotations Thus far we have been using all LLM annotations available. It is also intriguing to check the performance when we limit the volume of annotations, to simulate more resource-constrained settings. Here we subsample the LLM-annotated messages with rate 0.1 and 0.01 before sending them back to the hash table. Figure˜6 shows that the performance of ONI-retrieval significantly drops with rate 0.01, whereas ONI-classification remains comparable. This suggests that using a parametric reward model can help reduce the number of annotations required thanks to its generalization ability.
@@ A.2 Hyperparameters
Table˜1 shows the APPO hyperparameters which are common to all experiments and Table˜2 includes the hyperparameters for the online LLM annotation.
@@ Appendix B Prompts
B.1 Goal String and Prompts used in Section˜5.2
@@ B.1.2 Prompt used for ONI-retrieval
The prompt used for ONI methods have been modified to reflect the change of goals, which we list in the next subsection. In addition to the changes of goal description, we use the same examples as in Appendix˜B but changed its label.
@@ Appendix C Additional Results
In Figure˜7, we compare the performance of Motif using two different sized LLMs on Score and Oracle (LLaMA-3.1-8B-Instruct and LLaMA-3.1-70B-Instruct). Interestingly, we do not observe a significant difference between the two. This is in contrast to the previous work of (Klissarov et al., 2023), who found a significant difference between using LLaMA-2-70B-chat and LLaMA-2-7B-chat. This suggest that the smaller 8B model is sufficient for evaluating messages on these tasks, hence we use it in our experiments.
Figure 7: Performance of Motif using two different LLMs. Curves represent means and shaded regions represent standard errors over 55 seeds.
@@ C.2 Effect of the Episodic Term for ELLM-BoW
Our implementation of ELLM-BoW has included the episodic-count based normalization (6), which is key to the performance of ELLM-BoW. Figure˜8 shows that directly using ELLM-BoW as in Du et al. (2023a) ELLM-BoW, without the episodic term, failed to make progress in all three tasks.
@@ C.3 Additional Ablations
Impact of Intrinsic Reward Coefficient $\beta$ For ONI-classification and ONI-retrieval, we have used different values of $\beta$ for the Score task and other sparse reward tasks in Section˜5.1. Throughout our experiments, we have found that for these two methods, larger values of $\beta$ lead to better performance for the sparse reward tasks, while smaller values of $\beta$ better balance between intrinsic and extrinsic rewards for the dense reward Score task. In comparison, ONI-ranking is relatively robust to this choice, where larger values of $\beta$ are still slightly favored for the sparse reward tasks.
ONI-classification: the impact of the classification threshold As described in Section˜3.2, ONI-classification predicts binary rewards by modeling $P(y_{t}=1|o_{t})$ and then thresholding with $\eta$. Instead of using binary reward, an alternative design choice is to use real valued reward
@@
where $r^{\mathrm{int}}(o_{t})\in[0,1]$ is the output of the reward classifier before thresholding. Figure˜11 shows that the performance of ONI-classification on the four NetHack tasks is relatively robust to this hyperparameter $\eta$, and using $P(y_{t}=1|o_{t})$ as the reward leads to similar performance. As the training progresses, our reward model outputs values close to 0 or 1777To increase the diversity of the captions in the training dataset, we do not requery the LLM server if a caption is already annotated before. Therefore, every caption in our dataset only has a single label, resulting in this phenomenon., and we hypothesize this is the reason why the final performance remains comparable. Moreover, the natural classification threshold $\eta=0.5$ marginally outperforms the other values in the reward-free setting.
@@
Figure˜12 shows the performance and the number of captions stored in the replay buffer for both schemes. We see that for the deduplicated variant, the number of captions does not grow past a certain point (approximately $70k$ unique captions). For the non-deduplicated variant, the number of captions keeps growing linearly over time. The deduplicated variant fails to learn, which highlights the important effect which the annotation dataset can have. We hypothesize that the deduplicated variant may undersample captions that occur frequently in the agent’s experience, for which it is important to reliably estimate reward. For example, the blank message occurs very frequently during policy learning, but is only included in a small fraction of the pairs sent to the LLM for annotation, since it is sampled with the same probability as the other $\sim 70k$ other captions. In contrast, the duplicated variant samples the blank message and other frequent captions with much higher probability. Still, this remains a simple strategy, and designing more sophisticated sampling mechanisms (for example, that account for the epistemic uncertainty of the reward model) would be an interesting direction for future work.